A ball rolls off a 1.10-meter high table with an initial horizontal velocity of 3.3 m/s. Determine the time required for the ball to fall to the ground and the horizontal distance between the table's edge and the ball's landing location
1
Expert's answer
2020-11-19T09:00:09-0500

We can find the time required for the ball to fall to the ground from the equation of motion of the ball in vertical direction:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}},""t=\sqrt{\dfrac{2\cdot 1.10\ m}{9.8\ \dfrac{m}{s^2}}}=0.47\ s."

We can find the horizontal distance between the table's edge and the ball's landing location from the equation of motion of the ball in horizontal direction:


"x=v_0t=3.3\ \dfrac{m}{s}\cdot 0.47\ s=1.55\ m."

Answer:

"t=0.47\ s."

"x=1.55\ m."